Solve
x²+(x+2)²=34
All I have so far is
x²+(x+2)²-34=0

Answer by ewatrrr(24785)

 
Hi,
Solve
x²+(x+2)²=34
All I have so far is
x²+(x+2)²-34=0 || Important. Good work.
x^2 + x^2 + 4x + 4 - 34 = 0 Expand, factor , solve: x = 3 0r x = -5
2x^2 + 4x - 30 = 0
(2x -6)(x+5)= 0